# Privacy-Aware Detection of Fake Identity Documents: Methodology, Benchmark, and Improved Detection Methods (FakeIDet2)

### v1

@@ 5.3 Evaluation Metrics
Similar to previous approaches presented in the literature [33], we adopt the ISO/IEC 30107-3 standard [^9] https://www.iso.org/standard/79520.html for the evaluation of fake ID detection technology. This includes the use of the Bona-fide Presentation Classification Error Rate (BPCER) and Attack Presentation Classification Error Rate (APCER) metrics. The BPCER (Bona Fide Presentation Classification Error Rate), Eq. 11, where $N_{BF}$ is the number of bona fide samples, $RES_{i}$ is a term that is equal to 1 if the presented sample is incorrectly classified as an attack, and $\tau$ is the operational point, which is a mere threshold used to classify a sample as bona fide if the confidence of the prediction goes below and as an attack if it goes over said value. These measures the proportion of bona fide samples (i.e., real IDs) that are incorrectly classified as attacks (i.e., fake IDs).
$BPCER(\tau)=\frac{\sum^{N_{BF}}_{i=1}RES_{i}}{N_{BF}}$ (11)
Conversely, the APCER (Attack Presentation Classification Error Rate), Eq. 12, quantifies the proportion of attack samples (i.e., fake IDs) that are misclassified as bona fide (i.e., real IDs).
$APCER(\tau)=\frac{1}{N_{BF}}\sum^{N_{BF}}_{i=1}1-RES_{i}$ (12)
@@ 6 Experimental Results
This section evaluates the performance of our proposed FakeIDet2. The performance results included in all tables are at ID level, not patch level, over the final evaluation datasets. This facilitates the comparison of our proposed FakeIDet2 with traditional approaches in the literature, focused on the whole ID level evaluation.
First, we evaluate the technical novelties of our proposal. Sec. 6.1 provides a comprehensive comparison in terms of the softmax-based loss functions. The analysis of the optimal patch size configuration and anonymization level is carried out in Sec. 6.2 and Sec. 6.3, respectively. Finally, for completeness, we compare FakeIDet2 with our previous FakeIDet [23]. These first experiments are carried out using our novel FakeIDet2-db, as described in Sec. 5.2.
Once the preliminary analysis is conducted using our novel FakeIDet2-db, we select the optimal configuration balancing performance and privacy and use it to evaluate FakeIDet2 in two common scenarios for PAD. The first scenario is introduced in Sec. 6.4, which studies the performance of FakeIDet2 to unseen attacks. Hence, we train FakeIDet2 with the optimal configuration leaving one of the PAIs out of the training set. The second scenario, Sec. 6.5, analyzes the performance of FakeIDet2 when it is exposed to images taken with smartphones cameras which has not seen in training. To do so, we train FakeIDet2 using images from all camera sensors except one. Finally, we evaluate in Sec. 6.6 the best configuration of FakeIDet2 in a cross-database scenario, providing a standard public benchmark to the research community.
@@ 6.1 Selecting the Optimal Loss Function Configuration
  • Static-Weights: Each class has its own weight based on their frequency in the dataset (Eq. 5), and it is kept fixed along the training.
Table 4 shows the results for the different loss functions, considering our proposed dynamic class weights. We observe that for all loss functions, Dynamic-Weights achieves better overall results. For example, in the case of CosFace we see that the No-Weights and Static-Weights configurations achieve practically the same results with 4.32% EER, while the Dynamic-Weights improves a bit the performance, with 3.66% EER. ArcFace shows minor improvements using Dynamic-Weights, with an improvement of 0.34% EER and 1.67% EER with respect to using Static-Weights and No-Weights, respectively. The best performance is obtained with AdaFace using Dynamic-Weights, with 2.01% EER, providing an improvement of 1.65% EER over No-Weights, and 4.30% EER over Static-Weights.
@@
Table 4: Results in terms of EER (%) for the different loss functions (CosFace, ArcFace and AdaFace) through our proposed dynamic class weights (Dynamic-Weights). Dynamic-Weights allows to deal with class imbalance among the different PAIs patches. Results are achieved using the final evaluation of FakeIDet2-db, considering the configuration $\times128$ Non-Anonymized.
@@
Additionally, in order to prove that the improvement in performance correlates with how the embedded space is modeled, we plot in Fig. 7 the embeddings distributions of AdaFace using t-SNE projection into 2D. In the figure, we show visual examples when: i) No-Weights are applied (left), ii) Static-Weights are applied (middle), and iii) Dynamic-Weights are applied (right). The data used for this visualization was randomly sampled from the test set, comprising 6,000 patches.
In the No-Weights case, we observe that physical composite patches (red points) overlap with real patches (blue points). This is a critical issue, as confusion between these two classes can significantly hinder the detection of physical composite attacks at the document level—indicating that the embeddings are not sufficiently discriminative.
For Static-Weights, we observe a clearer clustering trend for physical composite patches. However, as expected, keeping the weight values fixed throughout training biases the model away from properly clustering patches from the other classes (i.e., real, print, and screen), which appear more scattered across the embedding space.
In contrast, with our proposed Dynamic-Weights, we observe a clear separation between physical composite patches and those from the remaining classes. This is enabled by the initial class weights computed at the start of training. Moreover, the clustering of the other classes is more consistent, with less sparsity in the outer sections of the space and more sparsity in the center compared to No-Weights and Static-Weights, which indicates a more robust feature extraction. We attribute this to the later training stages, during which the adaptive weights guide the model to better cluster patches from the more prevalent classes in the dataset.
The results of this section showcase that our proposed dynamic class weights approach suits better the problematic scenario of heavy class-imbalance. In summary, we adopt AdaFace loss function with Dynamic-Weights for the rest of our experiments.
@@
Table 5: Performance results for the fine-tuned DINOv2 (whole ID), FakeIDet [23], and proposed FakeIDet2 in terms of EER. Results are achieved using the evaluation set of FakeIDet2-db. The Non-Anonymized configuration is selected to simulate current scenarios in literature.
For this section, we consider the Non-Anonymized configuration in order to provide a direct comparison of our proposed FakeIDet2 approach, based on patches, with the current scenario in the literature, based on the whole Non-Anonymized ID. For completeness, we also include in the comparison our previous FakeIDet detector [23], trained using $\times128$ and $\times64$ patch sizes, similar to FakeIDet2. For the scenario of feeding the fake detector with the whole ID, we fine-tune DINOv2 [26] foundation model using a final sigmoid layer attached DINOv2’s output embedding.
Table 5 shows a comparison between the performance of the fine-tuned DINOv2 (whole ID), FakeIDet [23], and FakeIDet2 in terms of EER. Overall, we see that training on the whole ID does not achieve good results (22.63%), in contrast to our proposed patch-wise approach.
In the $\times128$ configuration, we observe that FakeIDet2 performs better than FakeIDet at detecting fake IDs of type screen (1.23% vs. 7.10% EER) and composite (2.09% vs. 54.08% EER), and remains close to FakeIDet in detecting print fake IDs (0.12% vs. 0.00% EER). Additionally, we observe that for the physical composite fake IDs, FakeIDet performance drops to 54.08% EER, while FakeIDet2 accurately detects these types of fake IDs leveraging the patch feature extraction and fusion, achieving a 2.09% EER. The drop in terms of performance is attributed to how FakeIDet computes the document level score. Given all the patches from a real or fake ID, it computes a score per individual patch and fuse them all together using a simple average, that yields a final score at document level. FakeIDet is reliable if the considered PAIs have the same patterns in the whole image. For example, in the case of a screen or print fake ID, it has the same properties over the whole ID (i.e., traces or fingerprints from the fake ID), but this is not true for physical composite fake IDs, as the ID is mostly real with the exception of the altered sections, failing to detect these kind of fake IDs properly. Finally, in the $\times64$ configuration, we can observe a similar trend, although performance seems to drop a bit in both FakeIDet and FakeIDet2.
These results prove the technical contributions of FakeIDet2 over FakeIDet [23], i.e., Patch Embedding Extractor and the Patch Embedding Fusion, that not only allows to detect attacks where the whole ID is a fake (i.e., print or screen), but also attacks where small sections are fake (i.e., composite). Given that the performance using the $\times64$ patch size configuration is close to the $\times128$ patch size, we select this patch size for further experiments, as it increases privacy in real-world settings, which is the main purpose of the study.
@@
In this section we analyze the impact in performance when removing sensitive information from an ID. To do so, we train FakeIDet2 in the $\times64$ patch size configuration at all anonymization levels, and evaluate the performance using Non-Anonymized data. This scenario simulates realistic real-world settings in which ID Holders have some restrictions when sharing IDs to AI Researchers in order to train the fake ID detectors. Table 6 shows the performance of FakeIDet2 in terms of EER (%) for the different anonymization levels and attacks. A particular trend can be observed, as performance per attack and overall shows that the EER increases as the amount of sensitive information decreases, which shows that FakeIDet2 performance depends on the amount of sensitive information available in the training stage (3.99% in Non-Anonymized vs. 8.64% in Pseudo-Anonymized vs. 17.94% in Fully-Anonymized). Even with this drop in terms of performance, we observe that the Pseudo-Anonymized $\times64$ configuration performs fairly well for only having seen partial information of the ID, which support our proposed hypothesis that training a model not using all the sensitive information of the ID produces quite competitive results and a small loss in performance as a trade-off. Given that this preliminary analysis confirmed that training with Pseudo-Anonymized data using $\times64$ patch size remains competitive, we decided to use this configuration for the rest of our experiments as it is balanced in terms of privacy and performance, which is the purpose of the study.
@@ 6.4 Leave-One-Attack-Out Scenario
Table 7 provides the results of FakeIDet2 when leaving different attacks out of the training data. For completeness, we include the EER for FakeIDet2 when trained using all PAIs available (All-Attacks). Based on our previous findings, we consider for training FakeIDet2 the Pseudo-Anonymized $\times64$ patch configuration, due to privacy reasons. The final evaluation is carried using Non-Anonymized $\times64$ patch configuration.
The worst results in terms of EER are obtained when composite attacks are left out, with a 28.24% when evaluating across all types of attacks. This shows that the physical composite attacks included in our novel FakeIDet2-db are the harder attacks for our model to detect, with 50.96% EER. It is also interesting to highlight that our screen attacks seem difficult to detect when they are out of training, with 36.00% EER. Finally, in the case of the print attacks, FakeIDet2 performs much better compared to detecting the other attacks with just a 6.77% in terms of EER when only detecting print attacks, although a little caveat should be considered. As covered in Sec. 3, our physical composite attacks are created by using small crops from print attacks, and lying them over real IDs. We argue that FakeIDet2 may be leveraging those learned features from composite patches to predict print attacks appropriately, giving competitive results that even improve the perfomance when the whole dataset is available (8.64% EER vs. 7.97% EER).
Finally, an interesting phenomena is observed when training FakeIDet2 leaving one attack out: the fake detector seems to “specialize” in detecting the attacks that are used in training, since the EER becomes even lower compared to the case of training using all attacks. For example, when leaving the composite attacks, FakeIDet2 obtains an EER of 2.02% when detecting print attacks, compared to 5.01% which is obtained when trained with all attacks. This can be expected, as when the number of classes is smaller, the data distribution to model by FakeIDet2 becomes simpler, hance becoming better at detecting the only attacks presented at training time.
@@ 6.5 Leave-One-Sensor-Out Scenario
Table 8 provides the results of FakeIDet2 when leaving different acquisition sensors out of the training data. For completeness, we include the EER for FakeIDet2 when trained using all devices available (All-Devices). Based on our previous findings, we consider for training FakeIDet2 the Pseudo-Anonymized $\times64$ patch configuration, due to privacy reasons. The final evaluation is carried using Non-Anonymized $\times64$ patch configuration.
As can be seen in the table, performance drops when the sensor with the highest quality (i.e., iPhone 15) is not used in training but only in the final evaluation. This seems not to be happening in the rest of scenarios, where images from the iPhone 15 are present in the training set. Given the different experiments that we run regarding this issue, our intuition on why this is happening relies on the fact that iPhone 15 cameras capture finer details that give more distinct features compared to Xiaomi’s and Redmi’s sensors. These fine grain characteristics allow for our FakeIDet2 to focus on traces that are more subtle, such as Moiré patterns on the screen attacks and the droplets of ink from the print and composite attacks, given that these attacks were printed with a Inkjet-kind printer. Furthermore, one of the main features of AdaFace loss function is that the adaptive margins change depending on the image quality. When using images of high-quality along with low- or medium-quality, the database is more diverse, and AdaFace helps the learning process by establishing larger margins to high-quality images and lower to low-quality. If images of low-quality are the majority of the database, the margins may not be as large as desired for proper separation between classes, and the features per each class may be overlapped in the embedded space, which are not helpful when forwarded into the Patch Embedding Fusion module, harming the performance. Evidences of this is the fact that results when images from the iPhone 15 camera are present, the evaluation performance of FakeIDet2 seems to be similar in both cases (i.e., No-Xiaomi vs. No-Redmi vs. All-Devices).

### v2

@@ 5.3 Evaluation Metrics
Similar to previous approaches presented in the literature [34], we adopt the ISO/IEC 30107-3 standard [^10] https://www.iso.org/standard/79520.html for the evaluation of fake ID detection technology. This includes the use of the Bona-fide Presentation Classification Error Rate (BPCER) and Attack Presentation Classification Error Rate (APCER) metrics. The BPCER measures the proportion of bona fide samples (i.e., real IDs) that are incorrectly classified as attacks (i.e., fake IDs), see Eq. 11, where $N_{\textrm{BF}}$ is the number of bona fide samples, $\textrm{RES}_{i}$ is a term that is equal to 1 if the presented sample is incorrectly classified as an attack, and $\tau$ is the operational point, i.e., the threshold used to classify a sample as bona fide if the confidence of the prediction goes below and as an attack otherwise.
$\textrm{BPCER}(\tau)=\frac{\sum^{N_{\textrm{BF}}}_{i=1}\textrm{RES}_{i}}{N_{\textrm{BF}}}$ (11)
Similarly, APCER quantifies the proportion of attack samples (i.e., fake IDs) that are misclassified as bona fide (i.e., real IDs):
$\textrm{APCER}(\tau)=\frac{1}{N_{\textrm{BF}}}\sum^{N_{\textrm{BF}}}_{i=1}1-\textrm{RES}_{i}$ (12)
@@ 6 Experimental Results
This section evaluates the performance of our proposed FakeIDet2. The performance results included in all tables are at the ID level, not the patch level, on the final evaluation datasets. This facilitates the comparison of our proposed FakeIDet2 with traditional approaches in the literature, focused on the whole ID level evaluation.
First, we evaluate the technical novelties of our proposal. Sec. 6.1 provides a comprehensive comparison in terms of softmax-based loss functions. The analysis of the optimal patch size configuration and anonymization level is carried out in Sec. 6.2 and Sec. 6.3, respectively. Finally, for completeness, we compare FakeIDet2 with our previous FakeIDet [25]. These first experiments are carried out using our new FakeIDet2-db, as described in Sec. 5.2.
Once the preliminary analysis is conducted using FakeIDet2-db, we select the optimal configuration balancing performance and privacy and use it to evaluate FakeIDet2 in two common scenarios for PAD. The first scenario is introduced in Sec. 6.4, which studies the performance of FakeIDet2 to unseen attacks. Hence, we train FakeIDet2 with the optimal configuration leaving one of the PAIs out of the training set. The second scenario, Sec. 6.5, analyzes the performance of FakeIDet2 when exposed to images taken with smartphone cameras that have not been seen in training. To do so, we train FakeIDet2 using images from all camera sensors except one. Finally, we evaluate in Sec. 6.6 the best configuration of FakeIDet2 in a cross-database scenario, providing a standard public benchmark for the research community.
@@ 6.1 Selecting the Optimal Loss Function Configuration
  • Static-Weights: Each class has its own weight based on its frequency in the dataset (Eq. 5), which is kept fixed during training.
Table 4 shows the results for the different loss functions and different types of class weights during learning. We observe that for all loss functions, the proposed Dynamic-Weights achieve better overall results. For example, in the case of CosFace we see that the No-Weights and Static-Weights configurations achieve practically the same results with 4.32% EER, while the Dynamic-Weights improve performance to 3.66% EER. ArcFace shows minor improvements using Dynamic-Weights, with an improvement of 0.34% EER and 1.67% EER with respect to using Static-Weights and No-Weights, respectively. The best performance is obtained with AdaFace using Dynamic-Weights, with 2.01% EER, providing an improvement of 1.65% EER over No-Weights, and 4.30% EER over Static-Weights.
@@
Table 4: Results in terms of EER (%) for the different loss functions (CosFace, ArcFace, and AdaFace) and different types of class weights during learning. The proposed Dynamic-Weights allow to deal with class imbalance among the different PAIs patches. Experiments are conducted on the entire evaluation set of FakeIDet2-db, considering the configuration $\times128$ Non-Anonymized.
@@
Furthermore, in order to prove that the improvement in performance correlates with how the embedded space is modeled, we plot in Fig. 7 the embeddings distributions of AdaFace using the t-SNE projection in 2D. In the figure, we show visual examples when: i) No-Weights are applied (left), ii) Static-Weights are applied (middle), and iii) Dynamic-Weights are applied (right). The data used for this visualization were randomly sampled from the test set, comprising 6,000 patches.
In the No-Weights case, we observe that the physical composite patches (red points) overlap with the real patches (blue points). This is a critical issue, as confusion between these two classes can significantly hinder the detection of physical composite attacks at the document level, indicating that the embeddings are not sufficiently discriminative.
For Static-Weights, we observe a clearer clustering trend for physical composite patches. However, as expected, keeping the weight values fixed throughout the training biases the model away from properly clustering patches from the other classes (i.e., real, print, and screen), which appear more scattered across the embedding space.
In contrast, with our proposed Dynamic-Weights, we observe a clear separation between physical composite patches and those from the remaining classes. This is enabled by the initial class weights computed at the start of training. Moreover, the clustering of the other classes is more consistent, with less sparsity in the boundaries between clusters and more sparsity within clusters compared to No-Weights and Static-Weights, which indicates a more robust feature extraction. We attribute this behavior to the later training stages using adaptive weights, which guide the model to better cluster patches from the more prevalent classes in the dataset.
The results of this section showcase that our proposed dynamic class weights approach better suits the problematic scenario of heavy class-imbalance. In summary, we adopt the AdaFace loss function with Dynamic-Weights for the rest of our experiments.
@@
For this section, we consider the Non-Anonymized configuration in order to provide a direct comparison of our proposed FakeIDet2 approach, based on patches, with the current scenario in the literature, based on the entire Non-Anonymized ID. For completeness, we also include in the comparison our previous FakeIDet detector [25], trained using $\times128$ and $\times64$ patch sizes, similar to FakeIDet2. For the scenario of feeding the fake detector with the whole ID, we fine-tune the DINOv2 foundation model [27] using a final sigmoid layer attached to the DINOv2 output embedding.
Table 5 shows a comparison between the performance of the fine-tuned DINOv2 (whole ID), FakeIDet [25], and FakeIDet2 in terms of EER. In general, we see that training on the whole ID does not achieve good results (22.63% EER), in contrast to our proposed patch-wise approach.
@@
Table 5: Performance results for the fine-tuned DINOv2 (Entire ID), FakeIDet [25], and proposed FakeIDet2 in terms of EER. Results are achieved using the evaluation set of FakeIDet2-db. The Non-Anonymized configuration (Entire ID) simulates current scenarios in literature.
In the $\times128$ configuration, we observe that FakeIDet2 performs better than FakeIDet in detecting fake IDs of type screen (1.23% vs. 7.10% EER) and composite (2.09% vs. 54.08% EER), and remains close to FakeIDet in detecting print fake IDs (0.12% vs. 0.00% EER). Furthermore, we observe that for the physical composite fake IDs, the performance of FakeIDet drops to 54.08% EER, while FakeIDet2 accurately detects these types of fake IDs using patch feature extraction and fusion, achieving a 2.09% EER. The performance drop is attributed to the way FakeIDet computes the document level score. Given all patches from a real or fake ID, it calculates a score per individual patch and fuses them all together using a simple average, which yields a final score at the document level. FakeIDet is reliable if the PAIs considered have the same patterns in the whole image. For example, in the case of a screen or print fake ID, it has the same properties throughout the whole ID (that is, traces or fingerprints of the fake ID), but this is not true for physical composite fake IDs, as the ID is mostly real with the exception of altered sections, failing to detect these types of fake ID properly. Finally, in the $\times64$ configuration, we can observe a similar trend, although performance seems to drop a bit in both FakeIDet and FakeIDet2.
These results prove the technical contributions of FakeIDet2 over FakeIDet [25], i.e., Patch Embedding Extractor and the Patch Embedding Fusion, that not only allows detecting attacks where the entire ID is a fake (i.e., print or screen), but also attacks where small sections are fake (i.e., composite). Given that the performance using the $\times64$ patch size configuration is close to the $\times128$ patch size, we selected this patch size for further experiments, as it increases privacy in real-world settings, which is the main purpose of the study.
@@
In this section, we analyze the impact on performance when removing sensitive information from an ID. To do so, we train FakeIDet2 in the $\times64$ patch size configuration at all anonymization levels, and evaluate the performance using Non-Anonymized data. This scenario simulates realistic real-world settings in which ID Holders have some restrictions when sharing IDs to AI Researchers in order to train the fake ID detectors. Table 6 shows the performance of FakeIDet2 in terms of EER (%) for the different anonymization levels and attacks. A particular trend can be observed: the fake ID detection errors (per attack and overall) increase as we reduce the amount of sensitive information, which shows that FakeIDet2 performance depends on the amount of sensitive information available in the training stage (3.99% in Non-Anonymized vs. 8.64% in Pseudo-Anonymized vs. 17.94% in Fully-Anonymized). Even with this drop in terms of performance, we observe that the Pseudo-Anonymized $\times64$ configuration performs fairly well despite having seen only partial information of the ID, which supports our proposed hypothesis that training a model not using all the sensitive information of the ID may generate a small loss in performance as a trade-off. Given that this preliminary analysis confirmed that training with Pseudo-Anonymized data using $\times64$ patch size remains competitive, we decided to use this configuration for the rest of our experiments, as it is balanced in terms of privacy and performance, which is the purpose of the study.
@@ 6.4 Leave-One-Attack-Out Scenario
Table 7 provides the results of FakeIDet2 when different attacks are left out of the training data. For reference, we include EER for FakeIDet2 when trained using all PAIs available (All-Attacks). Based on our previous findings, we consider for training FakeIDet2 the Pseudo-Anonymized $\times64$ patch configuration, due to privacy reasons. The final evaluation is carried using Non-Anonymized $\times64$ patch configuration.
The worst results in terms of EER are obtained when composite attacks are left out, with a 28.24% when evaluating all types of attacks. This shows that the physical composite attacks included in our new FakeIDet2-db are the harder attacks for our model to detect, with 50.96% EER. It is also interesting to note that our screen attacks seem difficult to detect when they are out of training, with 36.00% EER. Finally, in the case of the print attacks, FakeIDet2 performs much better compared to detecting the other attacks with just a 6.77% in terms of EER when only detecting print attacks, although a little caveat should be considered. As covered in Sec. 3, our physical composite attacks are created by using small crops from print attacks and lying them over real IDs. We argue that FakeIDet2 may be leveraging those learned features from composite patches to predict print attacks appropriately, giving competitive results that even improve the performance when the whole dataset is available (8.64% EER vs 7.97% EER).
Finally, an interesting phenomenon is observed when training FakeIDet2 leaving one attack out: the fake detector seems to “specialize” in detecting the attacks that are used in training, since the EER becomes even lower compared to the case of training using all attacks. For example, when leaving the composite attacks, FakeIDet2 obtains an EER of 2.02% when detecting print attacks, compared to 5.01% which is obtained when trained with all attacks. This can be expected, as when the number of classes is smaller, the data distribution to model by FakeIDet2 becomes simpler, hence becoming better at detecting the only attacks presented at training time.
@@ 6.5 Leave-One-Sensor-Out Scenario
Table 8 provides the results of FakeIDet2 when different acquisition sensors are left out of the training data. For reference, we include EER for FakeIDet2 when trained using all available devices (All-Devices). Based on our previous findings, we consider for training FakeIDet2 the Pseudo-Anonymized $\times64$ patch configuration, due to privacy reasons. The final evaluation is carried using Non-Anonymized $\times64$ patch configuration.
As can be seen in the table, performance decreases when the highest quality sensor (i.e., iPhone 15) is not used in training but only in the final evaluation. This seems not to be happening in the rest of the scenarios, where images from the iPhone 15 are present in the training set. Given the different experiments that we run regarding this issue, our intuition on why this is happening relies on the fact that iPhone 15 cameras capture finer details that give more distinct features compared to Xiaomi’s and Redmi’s sensors. These fine-grain characteristics allow our FakeIDet2 to focus on traces that are more subtle, such as Moiré patterns on the screen attacks and the droplets of ink from the print and composite attacks, given that these attacks were printed with an Inkjet-kind printer. Furthermore, one of the main features of the AdaFace loss function is that the adaptive margins change depending on the image quality. When using images of high quality along with images of low or medium quality, the database is more diverse, and AdaFace helps the learning process by establishing larger margins for high quality images and lower for low quality images. If images of low quality are the majority of the database, the margins may not be as large as desired for proper separation between classes, and the features per each class may be overlapped in the embedded space, which are not helpful when forwarded into the Patch Embedding Fusion module, harming the performance.